OpenStudy (anonymous):

find the linear approximation L(x) of the function f(x)=cos(9x) at a=pi/2 is L(x)=A+Bx

OpenStudy (bahrom7893):

okay workin on it

OpenStudy (bahrom7893):

L(x) = f(a) + f'(a)(x - a)

OpenStudy (bahrom7893):

f'(x) = -Sin(9x) * 9

OpenStudy (bahrom7893):

f'(pi/2) = -9Sin(9pi/2) = -9*1 = -9

OpenStudy (bahrom7893):

And f(pi/2) = Cos(9pi/2) = 0

OpenStudy (anonymous):

ok whats A and what B

OpenStudy (bahrom7893):

L(x) = 0 - 9(x-pi/2) L(x) = A + Bx (<--was this right, i keep forgetting the format?)

OpenStudy (anonymous):

yes but i do not know how u got ur answer at all

OpenStudy (bahrom7893):

Oh okay so let me explain... L(x) = f(a) + f'(a)(x - a) <-Linear approximation formula

OpenStudy (anonymous):

i got that

OpenStudy (bahrom7893):

Now for that you are going to need: f(a), f'(a) and x-a

OpenStudy (bahrom7893):

f(a) = f(pi/2)

OpenStudy (bahrom7893):

f(pi/2)=cos(9[pi/2]) = 0

OpenStudy (anonymous):

i got that the cos (9pi/2)= a decimal

OpenStudy (bahrom7893):

what is ur calculator's mode in? I bet it's in degrees.. supposed to be in radians

OpenStudy (bahrom7893):

pi/2 is a radian measure, not degree measure..

OpenStudy (bahrom7893):

try again? change the mode and find Cos(9pi/2)

OpenStudy (anonymous):

i am in radians

OpenStudy (bahrom7893):

okay can u go to mode and type in the values u have in here?

OpenStudy (anonymous):

dont kno how

OpenStudy (bahrom7893):

what calculator are u using?

OpenStudy (anonymous):

ti 83

OpenStudy (bahrom7893):

okay press MODE

OpenStudy (bahrom7893):

now type whatever u see in that screen here

OpenStudy (bahrom7893):

MODE is right next to 2nd key

OpenStudy (anonymous):

i see stuff like normal Sci Eng FLoat 0123456789 Radia Degree Func PAr Pol Seq

OpenStudy (bahrom7893):

what are the values, wait nevermind, can u take a picture of that mode screen and upload it on tinypic.com ? no registration needed btw

OpenStudy (anonymous):

dont know how to do that either

OpenStudy (bahrom7893):

do u have a camera?

OpenStudy (bahrom7893):

wait nevermind, trust me Cos(9pi/2) is 0

OpenStudy (bahrom7893):

okay?

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

L(x) = f(a) + f'(a)(x - a) <-Linear approximation formula f(a) = f(pi/2) f(pi/2)=cos(9[pi/2]) = 0

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

f'(x) = [cos(9x)]' = USE CHAIN RULE = -Sin(9x) * 9 = -9Sin(9x)

OpenStudy (bahrom7893):

f'(pi/2) = -9 Sin(9pi/2) = -9 * 1 = -9

OpenStudy (bahrom7893):

L(x) = f(a) + f'(a)(x - a) <-Linear approximation formula f(a) = f(pi/2) f(pi/2)=cos(9[pi/2]) = 0 f'(pi/2) = -9 Sin(9pi/2) = -9 * 1 = -9 L(x) = f(a) + f'(a)(x - a) L(pi/2) = 0 + -9 (x - pi/2)

OpenStudy (bahrom7893):

L(pi/2) = -9x - 9pi/2

OpenStudy (bahrom7893):

L(pi/2) = -9x - 9pi/2 L(x) = A + Bx L(pi/2) = -9pi/2 - 9x

OpenStudy (bahrom7893):

A = -9pi/2 B = -9

OpenStudy (bahrom7893):

Check the answer

OpenStudy (bahrom7893):

carra u here? I made a mistake, -9 (x - pi/2) is not -9pi/2 - 9x, its 9pi/2 - 9x

OpenStudy (bahrom7893):

Double negative is a positive... -9 * -pi/2 is 9pi/2 not -9pi/2